OpenStudy (anonymous):

x^2+y^2=25, what is the value of d^2y/dx^2 at the point (4,3)? (how do i solve using implicit differentiation?)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

yes you do

OpenStudy (misty1212):

twice

OpenStudy (solomonzelman):

you don't need implicit.

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle y=\sqrt{25-x^2}}\)

OpenStudy (misty1212):

no, but it will certainly make life easier!

OpenStudy (solomonzelman):

that is your f, because the point (4,3) is on the top half of the circle.

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle y=\sqrt{25-x^2}}\) differentiate twice, and plug in x=3

OpenStudy (misty1212):

if you do it that way, first derivative is \[\frac{-x}{\sqrt{25-x^2}}\] but then you have to find the derivative of that one!

OpenStudy (solomonzelman):

well, product rule (denominator written as neg exponent), wouldn't be so hard.

OpenStudy (misty1212):

if you use implicit diff, first derivative is \[y'=-\frac{x}{y}\] easier for me to deal with i think

OpenStudy (anonymous):

using implicit differentiation, how would you get the 2nd derivative?